# references.bib

### v1

@@ Abstract
While humans intuitively manipulate garments and other textiles items swiftly and accurately, it is a significant challenge for robots. A factor crucial to the human performance is the ability to imagine, a priori, the intended result of the manipulation intents and hence develop predictions on the garment pose. This allows us to plan from highly obstructed states, adapt our plans as we collect more information and react swiftly to unforeseen circumstances. Robots, on the other hand, struggle to establish such intuitions and form tight links between plans and observations. This can be attributed in part to the high cost of obtaining densely labelled data for textile manipulation, both in quality and quantity. The problem of data collection is a long standing issue in data-based approaches to garment manipulation. Currently, the generation of high quality and labelled garment manipulation data is mainly attempted through advanced data capture procedures that create simplified state estimations from real-world observations. In this work, however, we propose to generate real-world observations from given object states. To achieve this, we present GARField (Garment Attached Radiance Field) a differentiable rendering architecture allowing data generation from simulated states stored as triangle meshes. Code is available on the project website.
@@ I INTRODUCTION
Clothes and textile objects are omnipresent in everyday life, but their manipulation is still a long-standing challenge in robotics. This hinders the development of domestic, assistive robots that could drastically improve the quality of life of the many people in areas such as: nursing services, hospital and elderly care.
@@
The critical challenges of garment manipulation stems from the non-conformity of these objects with fundamental assumptions in robotics such as rigidity, known dynamics models, and low dimensional state space [longhini2024unfoldingliteraturereviewrobotic]. In the absence of such assumptions, state or partial-state estimation becomes an extremely high-dimensional prediction problem that is posed in conditions of partial observability.
@@
To circumvent this problem, some methods have abandonned state-estimation and rely on task-specific extrinsic features. Wrinkles detection, for instance, has proven efficient for simple tasks such as cloth flattening [sun_heuristic_based], [seita2019deeptransferlearningpick]. These features can be homogeneously defined over the whole surface of the object as they are not dependent on specific object-attached points. This allows perception systems to locate them from purely geometrical information, such as point-clouds or depth images without occlusion-reasoning. Furthermore, by default, they can be generalised to a wide range of objects of varying appearances. However, the absence of a notion of object state prevents their use in more general manipulation problems where understanding the topology of the object is crucial, e.g. dressing.
@@
Others have relied on visual and visuo-spatial feature recognition by training deep neural network to encode RGB or RGB-D observations in a constrained latent space [yan_learning_2020], [hoque_visuospatial_2021]. By using a full view of the object as input, these methods could address more complex manipulations tasks. But, learning these accurate and compressed latent representations for complex deformable objects requires a quantity of manipulation data that is presently unavailable.
@@
Despite the technical limitations of inferring a relevant state representation from real-world observations, several works have proven, in simulation, the usefulness of extensive and reliable state information to achieve skills such as dressing or dynamic manipulation [clegg_learning_2018], [9981376], [10414107]. However, during training, these methods rely on information that is unavailable in real world scenarios, hence they are either non-transferable, or they rely on perception transfer from simulated to real-world observations. The generation of perceptually convincing images for simulation is constrained by: 1) most simulated environments used for policy learning offer non-photorealistic rendering capabilities. 2) manually designing assets for high-fidelity rendering of real-life objects is a costly process requiring extensive human work.
@@
This means that the generation of high-quality images of simulated data forms a bottleneck in the development of generalisable, data-driven, advanced manipulation skills that can be used for manipulating complex textile items. In this work we address this problem, by proposing GARField (Garment Attached Radiance Field), a novel, learnable rendering pipeline allowing re-posing deformable meshes for realistic visual data generation from simulation states. Our contribution is three-fold:
@@
  • A simple training strategy to learning robust geometric and visual features from a limited numbers of static views
@@ II-A Differentiable Rendering
NeRF [mildenhall2021nerf], with its impressive rendering accuracy has been a cornerstone in establishing differentiable rendering as a relevant tool for multi-view scene capture. In parallel, DeepSDF [Park2019DeepSDFLC] has demonstrated the ability of deep neural architectures to accurately store an object’s geometry by fitting over point clouds. By combining these features, NeuS [wang2021neus], constrained NeRF’s underlying geometry representation to fit the definition of the signed distance function, thus enabling high quality capture of both a scene’s geometry and appearance. While these methods were mainly focused on rendering, they made it possible to integrate into a human-interpretable Euclidean 3D space, information from multiple point of views. This allows gradient flow to differentiable scene-embedded functions.
Object-NeRF [yang2021objectnerf], learnt distinct and composable object rendering functions for scene editing. PAC-NeRF [Li2023PACNeRFPA], PPR [yang2023ppr] and NDR [Cai2022NDR] paired differentiable rendering with physics-inspired deformation models producing coherent reconstruction over ordered sequences of images.
Differentiable rendering has also made its way into the robotics community where it has been used for grasp detection in Grasp-NeRF [Dai2023GraspNeRF] and Dex-NeRF [IchnowskiAvigal2021DexNeRF]. But also as a 3D decoder in [Li20213DNS] for learning a predictive model for visuo-motor control during fluid manipulation.
Finally, a similar approach has been used recently for data generation in the bin picking context with COV-NeRF [mishra2024closing]. COV-NeRF introduces an object-composable NeRF where the geometrical and visual characteristics of pickable objects are learnt in a first stage. In a second stage, the objects are re-arranged randomly to produce new and automatically labelled visuo-spatial data for downstream perception tasks. Our approach is similar in the general structure of the generation pipeline, however, because of the deformable nature of garments and the complex interaction they can have with their environment we have relied on stronger shape priors through using triangle meshes. Thies creates a significant divergences between the methods.
@@ II-B Visual sim-to-real gap in cloth manipulation
While the gap between simulated data and real-world observation is a problem encountered in all simulation-based robotics, it poses an even greater challenge in textile manipulation as cloth simulators generally produce inaccurate results unless slow and compute-intensive methods, such as FEM, are used [blancomulero2024benchmarking]. The visual sim-to-real gap is substantial problem resulting in significant performance degradation when transferring from simulation to real-world data as exemplified by [8967827] or [wu_learning_2023]. The main problem underlying this gap is the challenge of state estimation in a crumpled configuration.
Some works have tried to address this problem using shape reconstruction [chi2021garmentnets], [MEDOR] from half-crumpled observations. They have demonstrated interesting performance for state estimation.
In [SCRACT], the authors addressed the problem from the data generation angle by producing a dataset of observations with high quality annotations. This is similar to the approach in this work, however, we attain the same goal by automatic generation of convincing observations from known state rather than state estimation from observations with known action sequences.
Outside of robotic manipulation realistic renderings of garments has been tackled in the virtual try-on community [bhatnagar2019mgn], [corona2021smplicit] with more recent works producing outstanding results [10147324]. Nonetheless, the challenge of generating realistic garment images in the context of try-ons is simplified by the human body constraining the garment shape in a low-occlusion configuration. To the best of our knowledge, our work is the first to present a learnt automatic data generation pipeline for garments in any state.
@@ III METHOD
Figure 2: GARField models the scene as a composition of signed distance and visual feature fields. The background field is defined in the scene’s global coordinates frame. The other fields are attached to objects’ meshes and can be re-posed. The mesh-attached coordinates system, projects query points in a coordinate system made up of the point’s distance to the mesh’s surface and coordinates of the surface-projection of the query point in a bespoke coordinate system built around Laplacian-based position embeddings and barycentric coordinates.
@@
During scene capture, the deformable object is set in a rigid state by stretching it over a solid shape template with minimal self-occlusion. In this configuration, a set of observations is collected, this limited dataset is then used to train our composable scene model.
During the scene rendering stage, we use a deformable object simulator to re-pose the object mesh in an arbitrary number of poses, we then use our model to render novel observations from our learnt visual features.
@@ III-A COMPOSABLE SCENE RENDERING
Similar to [wang2021neus], we sub-divide the problem to represent the scene geometry on the one hand, and, on the other hand, a colour function that attributes a colour to each point in the scene. More precisely, as shown in Fig. 2, we model the scene implicitly through three functions:
@@
In order to separate object-centric representations from the background scene, we further subdivide the scene geometry and latent features field by defining $\mathrm{S}_{o}:\mathcal{B}\longrightarrow\mathbb{R}$ and $\mathrm{S}_{w}:\mathcal{B}\longrightarrow\mathbb{R}$ the objects and background scene’s signed distance function and $\mathrm{V}_{o}:\mathcal{B}\longrightarrow\mathbb{R}^{k}$ and the corresponding feature fields: $\mathrm{V}_{w}:\mathcal{B}\longrightarrow\mathbb{R}^{k}$.
@@
$\begin{split}\forall p\in\mathcal{B},\quad\mathrm{S}(p)&=\min(\mathrm{S}_{w}(p),\mathrm{S}_{o}(p,s))\\ \mathrm{V}(p)&=\operatorname*{arg\,min}_{(\mathrm{S}_{w}(p),\mathrm{S}_{o}(p,s))}(\mathrm{V}_{w}(p),\mathrm{V}_{o}(p,s))\end{split}$ (1)
We parametrize $\mathrm{D}$, $\mathrm{V}_{w}$ and $\mathrm{S}_{w}$ with MLPs. $\mathrm{S}_{o}(p,s)$, $\mathrm{V}_{o}(p,s)$ follow a custom architecture described below.
Rendering of images from our scene model is done by ray shooting through $\mathcal{B}$. For each pixel, the corresponding ray is defined as the set $\{\mathbf{p}(j)=\mathbf{o}+j\mathbf{v}|j\geq 0\}$, where $\mathbf{o}$ is the camera centre, $\mathbf{v}$ is the unit direction of the ray, and $j$ is the distance from the camera centre. For each pixel of $\mathcal{I}_{s}^{c}$, the colour is then calculated by integrating along the ray:
@@ III-B MESH-BASED FIELDS
The key challenge in defining mesh-based fields is in defining a surface-attached coordinate system as garments generally present complex, non-euclidean geometries.
@@
In practice, we compute the signed distance function $sdf_{\mathcal{M}}$ thanks to the libigl implementation of [1407857]. However, we learn residual geometry that would not be correctly represented by the mesh, by approximating the true signed distance function $\mathrm{S}_{o}$ with a residual MLP.
The challenge then becomes defining a continuous coordinate system on the surface of the mesh. A naive approach could be to define a 2D Cartesian coordinate system attached to each triangle, or define a point’s position in a face with it’s distance to two or three of the face’s corners. However, these coordinate systems would not be over-edge continuous. We define over-edge continuity as; the continuity of a surface value at the edge between to connected triangles. A more popular technique, known as UV-mapping, is to devise a projection of the mesh surface to a 2D plan and attribute to each node its coordinates in this plane. However, because of the garments’ complex geometries, using such techniques produces maps with varying resolution levels over the surface, and with a break in continuity along chosen cut-edges.
@@
We can notice that this definition attributes to each node its own embedding. Additionally, given $\gamma$ a path moving over the surface from $ABC$ to $BDC$, and $P^{\prime}$ the intersection point between $\gamma$ and $BC$. The coordinates of $P^{\prime}$ approaching from $ABC$ are:
@@
As a result, this embedding preserves the notion of proximity between nodes. Additionally, it carries over meshes many of the intuitions of the more familiar sine-cosine embedding used in Euclidian spaces. For instance, it acts as a frequency separation operator on the mesh’s surface, allowing for faster learning of our feature fields. This behaviour can be observed in Fig. 3b, where the first coordinate displays two local optima, the second and third have three, and the fourth has four, thus increasing the amount of variation. Finally, LPE’s coordinates value are in the $[-1,1]$ interval, providing an adequately normalized input to downstream neural networks. During our training we use an embedding size of $k=256$.
@@ III-C TRAINING
Figure 4: (a) Qualitative results for reconstruction of training images. (b) Qualitative results for re-posed meshes rendering. Top left: color, bottom left: masks, right: depth.
@@
We train our architecture by gradient descent optimizing with the following loss function:
@@
With $(\alpha,\beta,\gamma,\delta)$ a set of hyper-parameters each balancing the influence of each loss term.
The colour loss $\mathcal{L}_{colour}$ is defined as a L2-loss. Similarly, the depth loss constrains all points from of the depth image $\mathcal{D}_{s}^{c}$ to have a 0 signed distance value with a L2-loss:
@@
The Eikonal loss is a common component in differentiable rendering which constrains the norm of the gradient of the signed distance function to tend towards one. This can intuitively be understood since, by definition, moving a distance of one unit in the direction of the gradient of the signed distance function increases the signed distance value by one.
@@
This loss term is deactivated after ten thousands training iteration by setting $\delta=0$ to allow $\mathrm{S}_{o}$ to learn residual real-world geometry that is not modelled by $\mathcal{M}$.
@@
View direction is a necessary input to the colour rendering of the scene as it allows our model to learn how the light from the static light source bounces on surfaces. However, as we use a small number of static cameras in our scene, the view direction can be used by the network to over-fit to each camera and bias its output to fit to each camera individually instead of producing a cohesive general scene representation. This can, for instance, lead to poor generalization capabilities for novel views or to learning a strongly view-dependent feature field. To mitigate this effect, we employ a view direction augmentation by randomly replacing the view direction with a random sample from $\mathcal{S}\mathcal{O}(3)$, during training. In our trainings, we used a probability of 0.3 for replacement.

### v2

@@ Abstract
While humans intuitively manipulate garments and other textile items swiftly and accurately, it is a significant challenge for robots. A factor crucial to human performance is the ability to imagine, a priori, the intended result of the manipulation intents and hence develop predictions on the garment pose. That ability allows us to plan from highly obstructed states, adapt our plans as we collect more information and react swiftly to unforeseen circumstances. Conversely, robots struggle to establish such intuitions and form tight links between plans and observations. We can partly attribute this to the high cost of obtaining densely labelled data for textile manipulation, both in quality and quantity. The problem of data collection is a long-standing issue in data-based approaches to garment manipulation. As of today, generating high-quality and labelled garment manipulation data is mainly attempted through advanced data capture procedures that create simplified state estimations from real-world observations. However, this work proposes a novel approach to the problem by generating real-world observations from object states. To achieve this, we present GARField (Garment Attached Radiance Field), the first differentiable rendering architecture, to our knowledge, for data generation from simulated states stored as triangle meshes. Code is available on the project website.
@@ I INTRODUCTION
Clothes and textile objects are omnipresent in everyday life. However, their manipulation is still a long-standing challenge in robotics, which hinders the development of domestic assistive robots that could drastically improve the quality of life of many people in areas such as nursing services, elderly care and hospital care.
@@
The critical challenges of garment manipulation stem from the non-conformity of these objects with fundamental assumptions in robotics such as rigidity, known dynamics models, and low dimensional state space [longhini2024unfoldingliteraturereviewrobotic]. In the absence of such assumptions, state or partial-state estimation becomes an extremely high-dimensional prediction problem posed in partial observability conditions.
@@
To circumvent this problem, some methods have abandoned state estimation and rely on task-specific extrinsic features. Wrinkles detection, for instance, has proven efficient for simple tasks such as cloth flattening [sun_heuristic_based], [seita2019deeptransferlearningpick]. These features can be homogeneously defined over the whole surface of the object as they are not dependent on specific object-attached points, which allows perception systems to locate them from purely geometrical information, such as point clouds or depth images, without occlusion reasoning. Furthermore, by default, they can be generalised to a wide range of objects of varying appearances. However, the absence of a notion of object state prevents their use in more general manipulation problems where understanding the object’s topology is crucial, e.g. dressing.
@@
Others have relied on visual and visuo-spatial feature recognition by training deep neural networks to encode RGB or RGB-D observations in a constrained latent space [yan_learning_2020], [hoque_visuospatial_2021]. By using a full view of the object as input, these methods could address more complex manipulation tasks. However, learning these accurate and compressed latent representations for complex deformable objects requires a quantity of manipulation data that is presently unavailable.
@@
Despite the technical limitations of inferring a relevant state representation from real-world observations, several works have proven, in simulation, the usefulness of extensive and reliable state information to achieve skills such as dressing or dynamic manipulation [clegg_learning_2018], [9981376], [10414107]. However, during training, these methods rely on information unavailable in real-world scenarios. Hence, they are either non-transferable or rely on perception transfer from simulated to real-world observations. The generation of perceptually convincing images for simulation is constrained by: 1) most simulated environments used for policy learning offer non-photorealistic rendering capabilities. 2) manually designing assets for high-fidelity rendering of real-life objects is a costly process requiring extensive human work.
@@
As a result, generating high-quality images of simulated data forms a bottleneck in developing generalisable, data-driven, advanced skills for complex textile items’ manipulation. In this work, we address this problem by proposing GARField (Garment Attached Radiance Field), a novel, learnable rendering pipeline that allows re-posting deformable meshes for realistic visual data generation from simulation states. Our contribution is three-fold:
@@
  • A simple training strategy for learning robust geometric and visual features from a limited number of static views
In this article, we will first present related works in II, then formally pose the problem addressed and our method in III, and finally present our experiments, results and a short ablation study in IV before concluding.
@@ II-A Differentiable Rendering
With its impressive rendering accuracy, NeRF [mildenhall2021nerf] has been a cornerstone in establishing differentiable rendering as a relevant tool for multi-view scene capture. In parallel, DeepSDF [Park2019DeepSDFLC] has demonstrated the ability of deep neural architectures to accurately store an object’s geometry by fitting over point clouds. By combining these features, NeuS [wang2021neus] constrained NeRF’s underlying geometry representation to fit the definition of the signed distance function, thus enabling high-quality capture of both a scene’s geometry and appearance. While these methods mainly focused on rendering, they made it possible to integrate information from multiple points of view into a human-interpretable Euclidean 3D space. This architecture allows gradient flow to differentiable scene-embedded functions.
Object-NeRF [yang2021objectnerf], learnt distinct and composable object rendering functions for scene editing. PAC-NeRF [Li2023PACNeRFPA], PPR [yang2023ppr], and NDR [Cai2022NDR] paired differentiable rendering with physics-inspired deformation models producing coherent reconstruction over ordered sequences of images.
Differentiable rendering has also made its way into the robotics community, where it has been used for grasp detection in Grasp-NeRF [Dai2023GraspNeRF] and Dex-NeRF [IchnowskiAvigal2021DexNeRF]. But also as a 3D decoder in [Li20213DNS] for learning a predictive model for visuomotor control during fluid manipulation.
Finally, a similar approach has been used recently for data generation in the bin-picking context with COV-NeRF [mishra2024closing]. COV-NeRF introduces an object-composable NeRF that learns pickable objects’ geometrical and visual characteristics. Later, the objects are artificially re-arranged randomly to produce new and automatically labelled visuospatial data for downstream perception tasks. Our approach is similar in the general structure of the generation pipeline; however, because of the deformable nature of garments and the complex interaction they can have with their environment, we have relied on stronger shape priors by using triangle meshes, implying significant divergence between the methods.
@@ II-B Visual sim-to-real gap in cloth manipulation
While the gap between simulated data and real-world observation is a problem encountered in all simulation-based robotics, it poses an even greater challenge in textile manipulation as cloth simulators generally produce inaccurate results unless slow and compute-intensive methods, such as FEM, are used [blancomulero2024benchmarking]. The visual sim-to-real gap is a substantial problem resulting in significant performance degradation when transferring from simulation to real-world data, as exemplified by [8967827] or [wu_learning_2023]. The main problem underlying this gap is the challenge of state estimation in a crumpled configuration.
Some works have tried to address this problem using shape reconstruction [chi2021garmentnets], [MEDOR] from half-crumpled observations. They have demonstrated encouraging performance for state estimation.
In [SCRACT], the authors addressed the problem from the data generation angle by producing a dataset of observations with high-quality annotations. This is similar to the approach in this work, but we attain the same goal by automatic generation of convincing observations from known states rather than state estimation from observations with known action sequences.
Outside of robotic manipulation, realistic renderings of garments have been tackled in the virtual try-on community [bhatnagar2019mgn], [corona2021smplicit] with more recent works producing outstanding results [10147324]. Nonetheless, the challenge of generating realistic garment images in the context of try-ons is simplified by the human body constraining the garment shape in a low-occlusion configuration. To our knowledge, our work is the first to present a learnt automatic data generation pipeline for garments in any state.
@@ III METHOD
Figure 2: GARField models the scene as a composition of signed distance and visual feature fields. The background field is defined in the scene’s global coordinates frame. The other fields are attached to objects’ meshes and can be re-posed. The mesh-attached coordinates system projects query points in a coordinate system made up of the point’s distance to the mesh’s surface and coordinates of the surface-projection of the query point in a bespoke coordinate system built around Laplacian-based position embeddings and barycentric coordinates.
@@
During scene capture, the deformable object is set in a rigid state by stretching it over a solid shape template with minimal self-occlusion. In this configuration, a small set of observations is collected. This limited dataset is then used to train our composable scene model.
During the scene rendering stage, we use a deformable object simulator to re-pose the object mesh in an arbitrary number of poses; we then use our model to render novel observations from our learnt visual features.
@@ III-A COMPOSABLE SCENE RENDERING
Similar to [wang2021neus], we sub-divide the problem to represent the scene geometry on the one hand and, on the other hand, a colour function that attributes a colour to each point in the scene. More precisely, as shown in Fig. 2, we model the scene implicitly through three functions:
@@
To separate object-centric representations from the background scene, we further subdivide the scene geometry and latent features field by defining $\mathrm{S}_{o}:\mathcal{B}\longrightarrow\mathbb{R}$ and $\mathrm{S}_{w}:\mathcal{B}\longrightarrow\mathbb{R}$ the objects and background scene’s signed distance function and $\mathrm{V}_{o}:\mathcal{B}\longrightarrow\mathbb{R}^{k}$ and the corresponding feature fields: $\mathrm{V}_{w}:\mathcal{B}\longrightarrow\mathbb{R}^{k}$.
@@
$\begin{split}\forall p\in\mathcal{B},\quad\mathrm{S}(p)&=\quad\min(\mathrm{S}_{w}(p),\mathrm{S}_{o}(p,s))\\ \mathrm{V}(p)&=\operatorname*{arg\,min}_{(\mathrm{S}_{w}(p),\mathrm{S}_{o}(p,s))}(\mathrm{V}_{w}(p),\mathrm{V}_{o}(p,s))\end{split}$ (1)
We parametrise $\mathrm{D}$ with an MLP of 4 layers of hidden dimension 256, and $\mathrm{V}_{w}$ and $\mathrm{S}_{w}$ with an MLP of 8 layers of hidden dimension 256 and a skip connection at layer 4 . $\mathrm{S}_{o}(p,s)$, $\mathrm{V}_{o}(p,s)$ follow a custom architecture described below.
Rendering images from our scene model is done by ray shooting through $\mathcal{B}$. For each pixel, the corresponding ray is defined as the set $\{\mathbf{p}(j)=\mathbf{o}+j\mathbf{v}|j\geq 0\}$, where $\mathbf{o}$ is the camera centre, $\mathbf{v}$ is the unit direction of the ray, and $j$ is the distance from the camera centre. For each pixel of $\mathcal{I}_{s}^{c}$, the colour is then calculated by integrating along the ray:
@@ III-B MESH-BASED FIELDS
The key challenge in defining mesh-based fields is defining a surface-attached coordinate system, as garments generally present complex, non-euclidean geometries.
@@
In practice, we compute the signed distance function $sdf_{\mathcal{M}}$ thanks to the libigl implementation of [1407857]. However, we learn residual geometry that the mesh would not correctly represent by approximating the true signed distance function $\mathrm{S}_{o}$ with a residual MLP of 8 layers with hidden dimension 768 and a skip connection at layer 4.
The challenge then becomes defining a continuous coordinate system on the surface of the mesh. A naive approach could be to define a 2D Cartesian coordinate system attached to each triangle or define a point’s position in a face with its distance to two or three of the face’s corners. However, these coordinate systems would not be over-edge continuous. We define over-edge continuity as the continuity of a surface value at the edge between two connected triangles. A more popular technique, known as UV mapping, is to devise a projection of the mesh surface to a 2D plane and attribute to each node its coordinates in this plane. However, because of the garments’ complex geometries, using such techniques produces maps with varying resolution levels over the surface and with a break in continuity along chosen cut-edges.
@@
We can notice that this definition attributes to each node its own embedding. Additionally, given $\gamma$, a path moving over the surface from $ABC$ to $BDC$, and $P^{\prime}$, the intersection point between $\gamma$ and $BC$. The coordinates of $P^{\prime}$ approaching from $ABC$ are:
@@
As a result, this embedding preserves the notion of proximity between nodes. Additionally, it carries over meshes many of the intuitions of the more familiar sine-cosine embedding used in Euclidian spaces. For instance, it acts as a frequency separation operator on the mesh’s surface, allowing for faster learning of our feature fields. This behaviour can be observed in Fig. 3b, where the first coordinate displays two local optima, the second and third have three, and the fourth has four, thus increasing the amount of variation. Finally, LPE’s coordinates values are in the $[-1,1]$ interval, providing an adequately normalised input to downstream neural networks. During our training, we used an embedding size of $k=256$.
@@ III-C TRAINING
Figure 4: (a) Qualitative results for reconstruction of training images. (b) Qualitative results for re-posed meshes rendering. Top left: colour, bottom left: masks, right: depth. While the mesh position is more rigid that and usual cotton sock would be, we can observe alignment of geometry and colour features. Additionally, the quality of depth images and masks is on par with training data.
@@
We train our architecture by gradient descent optimising with the following loss function:
@@
With $(\alpha,\beta,\gamma,\delta)$ a set of hyper-parameters, each balancing the influence of each loss term.
The colour loss $\mathcal{L}_{colour}$ is defined as an L1-loss. Similarly, the depth loss constrains all points from the depth image $\mathcal{D}_{s}^{c}$ to have a 0 signed distance value with an L1-loss:
@@
The Eikonal loss is a common component in differentiable rendering, which constrains the norm of the gradient of the signed distance function to tend towards one. This can intuitively be understood since, by definition, moving a distance of one unit in the direction of the gradient of the signed distance function increases the signed distance value by one.
@@
This loss term is deactivated after ten thousand training iterations by setting $\delta=0$ to allow $\mathrm{S}_{o}$ to learn residual real-world geometry that is not modelled by $\mathcal{M}$.
@@
View direction is a necessary input to the colour rendering of the scene as it allows our model to learn how the light from the static light source bounces on surfaces. However, as we use a small number of static cameras in our scene, the network can use the view direction to over-fit each camera and bias its output to fit each camera individually instead of producing a cohesive general scene representation. This can, for instance, lead to poor generalisation capabilities for novel views or to learning a strongly view-dependent feature field. To mitigate this effect, we employ a view direction augmentation by randomly replacing the view direction with a random sample from $\mathcal{S}\mathcal{O}(3)$, during training. In our training, we used a probability of 0.3 for replacement.